# Observational prospects of self-interacting scalar superradiance with next-generation gravitational-wave detectors

### v1

@@ C.2 MOA-2011
Figure 11: Parameter space reachable by current and next-generation detectors (enclosed by the solid curves; see legend) for signals from MOA-2011 assuming parameters as in Table 2. Cases are shown for an assumed BH age of $1.0\times 10^{4}$ yr (top row), $1.0\times 10^{6}$ yr (middle row), and $1.0\times 10^{8}$ yr (bottom row). The reachable regions for 211 annihilation, 322 annihilation, and $322\rightarrow 211$ transition signals are shown in the left, middle, and right columns, respectively. The dashed curves divide the parameter spaces into three regimes: (top) the strong self-interaction regime, (middle) the moderate self-interaction regime, and (bottom) the gravitational regime. The reachable regions for Einstein Telescope (enclosed by black curve) and Cosmic Explorer (enclosed by red curve) largely overlap.

### v2

@@ C.2 MOA-2011
Figure 11: Parameter space reachable by current and next-generation detectors (enclosed by the solid curves; see legend) for signals from MOA-2011 assuming parameters as in Table 2. Cases are shown for an assumed BH age of $1.0\times 10^{4}$ yr (top row), $1.0\times 10^{6}$ yr (middle row), and $1.0\times 10^{8}$ yr (bottom row). The reachable regions for 211 annihilation, 322 annihilation, and $322\rightarrow 211$ transition signals are shown in the left, middle, and right columns, respectively. The dashed curves divide the parameter spaces into three regimes: (top) the strong self-interaction regime, (middle) the moderate self-interaction regime, and (bottom) the gravitational regime. The reachable regions for Einstein Telescope (enclosed by black curve) and Cosmic Explorer (enclosed by red curve) largely overlap. For reference, the QCD axion’s mass to $1/f$ relation is shown by a dotted white line (in the bottom right corner of each plot).